ADNAN MUHAMMED

Grade 12,

Let z be a complex number of constant modulus such that z*z is purely imaginary then the number of possible values of z is?

Let z be a complex number of constant modulus such that z*z is purely imaginary then the number of possible values of z is?

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

let z=x+iy, x^2 + y^2 is constant. z*z=x^2 - y^2 +i2xy is purely imaginary implies that x=±y,x^2+y^2=2x^2=k,k>0, so there are 2 possible values of z